# Domain-aware priors enable vertical federated learning in data-scarce coral multi-omics Sam Victor1*

### v1

@@ Conclusion
We demonstrate, to our knowledge, the first successful vertical federated learning on ultra-high-dimensional small-sample coral multi-omics, achieving robust thermal stress classification (AUROC 0.833±0.030) where standard VFL setups perform sub-optimally. Our work proposes three empirically motivated design principles for enabling federated learning under extreme PN regimes:
Design Principle 1: Domain-aware priors appear necessary in this regime. Generic VFL exhibits gradient noise domination (AUROC0.5) when P/N6,000, treating features as anonymous tokens and allowing irrelevant high-dimensional gradients to overwhelm signal. Domain-aware feature selection via gradient saliency transforms the problem from statistically infeasible (P=90,579) to tractable (P=1,300), enabling convergent learning. While our specific biological weights (transcriptomics 1.5×, microbiome 0.5×) are coral-centric, the underlying design principle, incorporating structured domain knowledge as feature priors, may generalize to other data-scarce domains with established knowledge: rare diseases (GWAS with pathway priors), microbiomes (phylogenetic constraints), precision agriculture (spectral physics), climate science (process models).
Design Principle 2: Stability metrics supersede peak performance in small-sample regimes. The 7-fold variance reduction relative to LASER (SD 0.030 vs 0.215) represents the critical practical advantage, even beyond the substantial AUROC improvement (+0.233, 39% relative gain), when deployment requires reliability across random seeds, data splits, and cohort variations.
Design Principle 3: Interpretable feature selection mitigates gradient instability and enables validation. Gradient saliency features (heat shock proteins, oxidative stress enzymes, photosystem markers) provide dual benefits: biologically plausible features stabilize gradient flow (Jaccard=1.0 consistency), and interpretability validates that performance reflects true signal, not overfitting. Interpretability is not post-hoc explanation but a mechanism supporting convergence and domain acceptance. These principles enable worldwide coral institutions to collaboratively train ML models without data sharing, addressing regulatory, competitive, and ethical barriers to multi-laboratory coordination.
By demonstrating privacy-preserving federated learning on N=13 samples with P=90,579 features, previously considered infeasible, this work opens avenues for distributed collaborative science in ecology, rare disease research, and other domains where data scarcity intersects with privacy constraints. The convergence of explainable AI, domain knowledge, and federated architectures represents a paradigm shift from data-centric to knowledge-centric machine learning. As coral reefs face accelerating climate threats (70% mortality projected by 2050), methodological advances enabling rapid knowledge synthesis across the global research community may prove as critical as the biological discoveries they enable. Our framework demonstrates that federation under scarcity is achievable when domain expertise guides feature engineering, transforming VFL from a data-abundant scalability solution into a knowledge-centric collaboration tool for science’s most urgent challenges.
@@ Acknowledgments
The author acknowledges the Stephens et al. research group for making the Montipora capitata multi-omics dataset publicly available via Zenodo, enabling reproducible coral stress research. The author also thanks the developers of NVIDIA FLARE, PyTorch, and related open-source projects for maintaining critical machine learning infrastructure. The author thanks Swarn S. W. for proofreading the final version of the manuscript and providing editorial feedback. The author is affiliated with Sri Eshwar College of Engineering as a student and with a research-oriented startup working on multi-omics platforms (Rikoukei Labs FZ-LLC) as a collaborator. However, this research was conducted independently and received no institutional or corporate funding, data, or computational resources. The author declares no competing interests.

### v3

@@ Conclusion
We demonstrate, to our knowledge, the first successful vertical federated learning on ultra-high-dimensional small-sample coral multi-omics, achieving robust thermal stress classification (AUROC 0.776±0.039) where standard VFL setups perform sub-optimally. Our work proposes three empirically motivated design principles for enabling federated learning under extreme PN regimes:
Design Principle 1: Aggressive dimensionality reduction enables above-chance learning in extreme PN regimes, while domain-aware priors are specifically necessary for stable, deployment-ready performance—confirmed by ablation (2.3 variance reduction, CV 0.050 vs 0.110 under equal-weights conditions).
Design Principle 2: Stability metrics supersede peak performance in small-sample regimes; the equal-weights ablation makes this concrete: a method with statistically indistinguishable mean AUROC () but a per-seed range of 0.714–0.952 is operationally less trustworthy than one bounded to 0.714–0.833.
Design Principle 3: Gradient saliency-guided feature selection provides dual benefits—biologically plausible features stabilize gradient flow, and interpretable outputs validate that performance reflects genuine signal rather than overfitting artifacts.
An equal-weights ablation additionally reveals that proteomics carries the strongest intrinsic discriminative signal for thermal stress classification in M. capitata under unbiased conditions (dominant in 65/65 fold-seed combinations, gradient importance 20 higher than transcriptomics), suggesting that data-driven prior calibration may further improve federated learning performance in future work. By demonstrating privacy-preserving federated learning on N=13 samples with P=90,579 features, previously considered infeasible, this work opens avenues for distributed collaborative science in ecology, rare disease research, and other domains where data scarcity intersects with privacy constraints. The convergence of explainable AI, domain knowledge, and federated architectures represents a paradigm shift from data-centric to knowledge-centric machine learning. As coral reefs face accelerating climate threats (70% mortality projected by 2050), methodological advances enabling rapid knowledge synthesis across the global research community may prove as critical as the biological discoveries they enable. Our framework demonstrates that federation under scarcity is achievable when domain expertise guides feature engineering, transforming VFL from a data-abundant scalability solution into a knowledge-centric collaboration tool for science’s most urgent challenges.
@@ Acknowledgments
The author acknowledges the Stephens et al. research group for making the Montipora capitata multi-omics dataset publicly available via Zenodo, enabling reproducible coral stress research. The author also thanks the developers of NVIDIA FLARE, PyTorch, and related open-source projects for maintaining critical machine learning infrastructure. The author thanks Swarn S. W. for proofreading the final version of the manuscript and providing editorial feedback. The author used AI-assisted tools (Claude, Anthropic) for statistical methodology review, numerical consistency verification, and as a learning resource; all scientific content, manuscript text, and figures were produced by the author. The author thanks the creators of matplotlib and seaborn for making paper figure creation accessible. The author is affiliated with Sri Eshwar College of Engineering as a student and with a research-oriented startup working on multi-omics platforms (Rikoukei Labs FZ-LLC) as a collaborator. However, this research was conducted independently and received no institutional or corporate funding, data, or computational resources. The author declares no competing interests.